# Stochastic analysis of finite-temperature effects on cosmological parameters by artificial neural networks

### v1

@@ 2.1 Modification of CLASS
The $\Lambda$CDM model (see [9, 10] for reviews) assumes a spatially flat or nearly flat Universe, with a cosmological constant $\Lambda$ representing dark energy. It successfully explains the large-scale structure of the Universe, CMB observations, and galaxy formation. The model involves six primary parameters (or seven if the curvature density parameter $\Omega_{K}$ is included), which govern the composition and evolution of the Universe. These parameters are:
$h,\quad\omega_{b},\quad\omega_{cdm},\quad A_{s},\quad n_{s},\quad\tau_{reio}.$ (4)
@@
The CLASS code [13, 14] is a computational tool used to solve the Boltzmann equations and generate the power spectra of cosmological observables within the $\Lambda$CDM framework. Once the cosmological parameters are specified, CLASS produces the relevant observables accordingly. However, when considering QG effects at finite temperatures, modifications to this standard framework become necessary. These arise due to the temperature dependence of QG effects, which become relevant at high energies and lead to shifts in cosmological parameters. Among these, the most prominent is the cosmological constant $\Lambda$ (denoted as $\Lambda_{1}$ in the present context), which becomes temperature-dependent due to quantum effects[3, 4, 5]. [^4] These works extend earlier studies of foliation-based gravity quantization at finite temperature [21, 22, 23, 24, 25]. In particular, the gauge $K=0$ (or more generally $K=K_{0}$), where $K$ is the trace of the second fundamental form, plays a key role [24]. This approach parallels Witten’s recent treatment in [26].
Referring to [5] and prior studies on the 1PI effective action for the Standard Model coupled to gravity, we find that the expression for the Hubble parameter $H(t)$ is modified. The original expression, as adopted in [9], is:
$H=7.204\times 10^{-19}T^{\frac{3}{2}}\sqrt{1.523\times 10^{-5}T+\left(\frac{2.725}{T}\right)^{3}\Omega_{\Lambda_{1}}h^{2}+\Omega_{M}h^{2}}.$ (5)
(In [9], both the $\Omega_{\Lambda_{1}}$ and curvature density parameter are neglected for their small contributions.) The modified CLASS implementation [18], [^5] The code was obtained in collaboration with K.-T. Cho and M.-S. Lee. includes finite-temperature corrections, yielding the following form:
$H$ $=$ $\sqrt{1.523\times 10^{-5}T+\left(\frac{2.725}{T}\right)^{3}\Omega_{\Lambda_{1}}h^{2}+\Omega_{M}h^{2}+\frac{T}{2.725}\Omega_{\Lambda_{2}}h^{2}+\frac{2.725}{T}\Omega_{\Lambda_{3}}h^{2}}$ (6)
×$\times 7.204\times 10^{-19}T^{\frac{3}{2}}.$
@@
A natural question arises as to whether $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ could be absorbed into existing parameters such as the radiation density $\Omega_{r}$ and curvature density $\Omega_{K}$, respectively, given their similar scaling behaviors. However, this is not possible: the perturbations arising from finite-temperature QFT effects are of a fundamentally different nature from those associated with radiation or curvature. For instance, $\Omega_{K}$ appears at the classical level in the standard $\Lambda$CDM cosmology whereas $\Omega_{\Lambda_{3}}$ has a strictly quantum origin. Hence, $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ are independent and not reducible to existing terms. (Also, see below.) The fluctuations $\delta\rho_{\Lambda_{2}}$ and $\delta\rho_{\Lambda_{3}}$ are neglected due to their suppressed nature — being both quantum in origin and small variations. Thus, their influence is minimal, and $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ effectively appear only within the modified Hubble parameter $H$. In contrast, radiation $\Omega_{r}$ directly contributes to the fluctuation equations and plays a significant role in the evolution of perturbations. In this work, $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ are treated phenomenologically, [^6] Theoretically, these parameters can be computed from first principles via the renormalization procedure. However, finite renormalization freedom necessitates experimental inputs like the present one to fix their values. and their inclusion is justified by consistency with numerical results.
One might initially assume that finite-temperature quantum loop-induced contributions to $\Lambda$ especially those involving photons and neutrinos - are redundant, since these are already accounted for as radiation. However, this conflation is misleading. Radiation arises from physical, on-shell particles (e.g., photons, neutrinos) through, say, kinetic theory physics, whereas the contributions under consideration are due to virtual particles in loop corrections. The finite-temperature effects we consider are analogous to vacuum energy contributions induced by loop effects in QED photon and electron fields. Radiation density, on the other hand, pertains to the energy density of physical photons and neutrinos. Therefore, the finite-temperature effects we consider are not included in the conventional radiation density and must be accounted for separately.
2.2 Statistical methods
In this study, we employ ANNs to predict the distance between the 2018 Planck data [15] and data generated by the modified CLASS with input cosmological parameter values. Fig. 2 represents the schematic diagram of the ANNs with eight nodes in the input layer representing the eight cosmological parameters and four hidden layers, each with 16 nodes and the output layer with one node of the distance. Once the ANN is trained, we use gradient-based optimization to determine the set of covariates that minimize the predicted distance. We then explain the feature ablation strategy to discuss the relative importance of cosmological parameters in the predictions. Finally, we developed the ANN-based probabilistic optimization to find the minimum of the distance, considering the numerical measurement errors.
2.2.1 Artificial neural networks
Let $\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta})$ represent a deep neural network with $L$ layers that transforms an input from $\mathbb{R}^{d_{\text{in}}}$ into an output in $\mathbb{R}^{d_{\text{out}}}$, where $d_{\text{in}}$ and $d_{\text{out}}$ denote the dimensions of the input and output spaces, respectively. Suppose $m_{l}$ represents the number of neurons in the $l$-th hidden layer of the ANN, where $m_{0}=d_{\text{in}}$ and $m_{L}=d_{\text{out}}$ corresponds to the input and output layer sizes, respectively. Each layer $l$ $(l=1,\dots,L-1)$ is associated with a weight matrix $\mathbf{W}^{l}\in\mathbb{R}^{m_{l}\times m_{l-1}}$, whose element $W^{l}_{j,k}$ denotes the weight connecting neuron $j$ in layer $l$ to neuron $k$ in layer $l-1$. In addition, each layer uses a bias vector $\mathbf{b}^{l}\in\mathbb{R}^{m_{l}}$, where the $j$-th element, $b_{j}^{l}$, represents the bias term associated with the $j$-th neuron in layer $l$. Suppose $\boldsymbol{\theta}=\{\mathbf{W}^{1},\dots,\mathbf{W}^{L},\mathbf{b}^{1},\dots,\mathbf{b}^{L}\}$ denotes the set of all unknown parameters of the ANN. Let $\mathbf{M}^{l}$ denote the response from layer $l$ before activation; hence $\mathbf{M}^{l}$ is given by:
$\mathbf{M}^{l}=\mathbf{W}^{l}\cdot\mathcal{F}^{l-1}(\mathbf{x},\boldsymbol{\theta})+\mathbf{b}^{l},\quad\text{for }l=1,\dots,L-1.$ (7)
$\mathcal{F}^{l}(\mathbf{x},\boldsymbol{\theta})$ $=\sigma(\mathbf{M}^{l}),\quad l=1,\dots,L-1,$ (8)
$\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta})$ $=\mathbf{M}^{L}.$
$\mathcal{J}(\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta}))=\sum_{i=1}^{n}(f(\mathbf{x}_{i})-\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta}))^{2},$ (9)
$\frac{\partial\mathcal{J}(\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta}))}{\partial\mathbf{M}^{l}}=\left(\mathbf{W}^{l+1}\right)^{\top}\cdot\frac{\partial\mathcal{J}(\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta}))}{\partial\mathbf{M}^{l+1}}\odot\sigma^{\prime}(\mathbf{M}^{l}),$ (10)
$\frac{\partial\mathcal{J}(\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta}))}{\partial\mathbf{M}^{L}}=\frac{\partial\mathcal{J}(\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta}))}{\partial\mathcal{F}^{L}(\mathbf{x},\boldsymbol{\theta})}\odot\sigma^{\prime}(\mathbf{M}^{L}).$ (11)
2.2.2 Gradient-based feature optimization with ANN
$\nabla\mathcal{F}^{L}({\bf x},\widehat{\boldsymbol{\theta}}_{opt})=\frac{\partial\mathcal{F}^{L}({\bf x},\widehat{\boldsymbol{\theta}}_{opt})}{\partial{\bf x}}.$ (12)
${\bf x}^{(t+1)}={\bf x}^{(t)}-\alpha\nabla\mathcal{F}^{L}({\bf x}^{(t)},\widehat{\boldsymbol{\theta}}_{opt})$ (13)
${\bf x}^{(t+1)}={\bf x}^{(t)}-{\bf H}_{t}^{-1}\left(\mathcal{F}^{L}({\bf x}^{(t)},\widehat{\boldsymbol{\theta}}_{opt})\right)\nabla\mathcal{F}^{L}({\bf x}^{(t)},\widehat{\boldsymbol{\theta}}_{opt}),$ (14)
where ${\bf H}_{t}\left(\mathcal{F}^{L}({\bf x}^{(t)},\widehat{\boldsymbol{\theta}}_{opt})\right)$ represents an approximation of the Hessian matrix, which ensures a more accurate search direction compared to standard gradient descent. The algorithm is repeatedly alternated until the stopping rule of $||\nabla\mathcal{F}^{L}({\bf x},\widehat{\boldsymbol{\theta}}_{opt})||_{2}<\epsilon$ is satisfied, where $||\cdot||_{2}$ represents the $\ell_{2}$ norm.
2.2.3 Feature ablation method
Let $\{({\bf x}_{i},y_{i});i=1,\ldots,n\}$ and $\{({\bf x}_{j,{\text{t}est}},y_{j,{\text{t}est}});j=1,\ldots,m\}$ denote, respectively, the training and test data of sizes $n$ and $m$ from the representing the Planck distance response function $y=f({\bf x})$ where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Because the objective response function is not explicitly available, as described in section 2.2.1, we first predict the objective function using the ANN methods based on the training data. Let $\mathcal{F}^{L}({\bf x},\widehat{\boldsymbol{\theta}}_{opt})$ denote the ANN-based estimate of the objective function $f({\bf x})$. The prediction accuracy performance of the trained ANN estimate in the presence of all the input features is computed based on MSE using test data [28]. Accordingly, the ANN-based MSE is computed by
$\text{MSE}\left(\mathcal{F}^{L}({\bf x},\widehat{\boldsymbol{\theta}}_{opt})\right)=\frac{1}{m}\sum_{j=1}^{m}\left(y_{j,{\text{t}est}}-\mathcal{F}^{L}({\bf x}_{j,{\text{t}est}},\widehat{\boldsymbol{\theta}}_{opt})\right)^{2}.$ (15)
Henceforth, the MSE (15) is called the full model MSE, showing the prediction error from an ANN-based estimate of the objective function using all the input features.
In the next step, we iteratively remove the individual feature $x_{j},j=1,\ldots,d$ from the input design matrix. Let ${\bf x}_{-j}$ denote the input features when the $j$-th feature $x_{j}$ has been removed from the input data. We then train the ANN estimator of the objective function similar to the full model and find the ANN-based estimate $\widehat{f({\bf x})}=\mathcal{F}^{L}({\bf x}_{-j},\widehat{\boldsymbol{\theta}}_{opt})$. Accordingly, we compute the reduced model MSE in a similar vein to equation (15) using the test data in the absence of the $x_{j}$ coordinates. Let $\text{MSE}\left(\mathcal{F}^{L}({\bf x}_{-j},\widehat{\boldsymbol{\theta}}_{opt})\right)$ denote the test MSE of the reduced model using data after removing the $j$-th coordinate for $j=1,\ldots,d$.
We then compute the feature ablation statistic after removing $x_{j}$ by
$\Delta\text{MSE}_{j}=\text{MSE}\left(\mathcal{F}^{L}({\bf x}_{-j},\widehat{\boldsymbol{\theta}}_{opt})\right)-\text{MSE}\left(\mathcal{F}^{L}({\bf x},\widehat{\boldsymbol{\theta}}_{opt})\right),$ (16)
2.2.4 ANN-based simulated annealing
${\bf x}^{*}={\bf x}^{(k)}+\Delta{\bf x},$
The new candidate ${\bf x}^{*}$ is then evaluated through a stochastic accept-reject method to see if ${\bf x}^{*}$ improves the minimum of the objective function or not. Since the objective function $f({\bf x})$ is unknown, the accept reject is applied to the ANN estimate of the objective function $\widehat{f}({\bf x})=\mathcal{F}^{L}({\bf x},\widehat{\boldsymbol{\theta}}_{opt})$ computed by the method described in section 2.2.1. Since the closed form of the ANN estimate is not available, the improvement of the new candidate ${\bf x}^{*}$ should be evaluated numerically. Hence, the cost difference based on ANN is computed by
$\Delta\mathcal{F}({\bf x}^{*};{\bf x}^{(k)})=\mathcal{F}^{L}({\bf x}^{*},\widehat{\boldsymbol{\theta}}_{opt})-\mathcal{F}^{L}({\bf x}^{(k)},\widehat{\boldsymbol{\theta}}_{opt}).$ (17)
$P(\text{accepting the transition from ${\bf x}^{(k)}$ to ${\bf x}^{*}$})=\begin{cases}1,&\text{if }\Delta\mathcal{F}({\bf x}^{*};{\bf x}^{(k)})\leq 0,\\ e^{-\Delta\mathcal{F}({\bf x}^{*};{\bf x}^{(k)})/T_{k}},&\text{if }\Delta\mathcal{F}({\bf x}^{*};{\bf x}^{(k)})>0,\end{cases}$ (18)
$T_{k+1}=\alpha T_{k},$ (19)
where $10<\alpha<$ 1 is the cooling rate. It is easy to see that in the first iteration of the SA, the high values of the temperature parameter enable the algorithm to explore a wider area of the parameter space to search better for a minimum. If the temperature drops too quickly, the algorithm may not be able to search the entire parameter scape and hence may converge to a suboptimal solution. Conversely, if the temperature is cooled too slowly, the algorithm requires many iterations to go from one area to another of the parameter space. Accordingly, the number of iterations increases, and the $\alpha$ rate controls the cooling temperature for balance exploitation and refinement of the minimum in the parameter space. The algorithm is alternated until no significant improvement is observed in the successive iterations or the algorithm reaches a pre-specified maximum number of iterations.
@@ 3 Numerical analysis
Our analysis employs a modified version of CLASS incorporating the quantum gravity modifications previously described. The parameter space exploration proceeds in two distinct phases. First, we conduct a comprehensive brute-force scan across eight cosmological parameters spanning generous ranges to ensure complete coverage of the parameter space. By these comprehensive scans tens of millions of data points are obtained. Following this initial broad exploration, we implement a refined search strategy focusing on regions that minimize the Euclidean distance between our computed power spectra and the Planck dataset. This targeted approach yields significantly improved precision, with the minimal achieved distance of 26.83 corresponding to parameter values largely consistent with existing literature. [^7] To give a sense of the accuracy associated with this number, the distance between the 2018 Planck curve and one generated by CLASS using the default.ini settings is approximately 75. Any curve with a distance to the Planck curve of less than approximately 1000 is visually indistinguishable from it. Notably, the optimal solution features a negative value for $\Omega_{\Lambda_{2}}$, a result that finds natural explanation through dimensional regularization in the renormalization procedure. The analysis also reveals an enhanced contribution from $\Omega_{\Lambda_{3}}$, suggesting potential degeneracy between finite-temperature effects and spatial curvature that warrants careful interpretation.
@@
To complement and enhance the efficiency of our parameter space exploration, we implement advanced ANN techniques. Various numerical calculations and machine learning methods have been recently applied to the physics of black hole solutions[40, 41, 42, 43, 37]. These machine learning methods enable precise identification of optimal parameter coordinates while maintaining computational tractability across the challenging eight-dimensional parameter space.
@@
Let us explore the estimation of the distance based on the seven (including nonzero $\Omega_{K}$) currently established parameters as well as the finite-T induced ones, $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$. [^8] In the 8-variable model $\Omega_{K}$ is set to zero. Throughout this section, the prediction accuracies are translated into the performance of the ANN-based method in detecting the minimum distance to the 2018 Planck curve from statistical populations in the presence or absence of the finite-T effects.
@@
Figure 3: Sampling distribution of the distance values based on eight cosmological parameters.
@@
The numerical analyses evaluate the distance between the 2018 Planck curve and 7-parameter and 8-parameter ones using 459,912 and 335,985 data points. These data points are computed over equally spaced grids spanning the domain of the two approaches. These two data sets are considered as two statistical populations corresponding to the two estimation methods. We consider the response variable $y$, representing the distance function $y=f({\bf x})$, which measures the deviation of the proposed estimation methods from the 2018 Planck power spectrum. In this setup, the parameters of the estimation methods were treated as explanatory variables ${\bf x}=(x_{1},\ldots,x_{J})$, where $J=7$ for the 7-parameter model and $J=8$ for the 8-parameter model. Using this framework, the ML model maps the cosmological parameters to the distance function, enabling us to evaluate and predict the performance of the two estimation methods. This approach was particularly valuable for assessing the accuracy of the proposed methods, in the presence or absence of $\Omega_{\Lambda_{2}}$.
@@
Figure 4: Sampling distribution of the distance values based on seven cosmological parameters.
@@
Figs. 3 and 4 show the sampling distributions of the distance responses based on two statistical populations derived from the entire dataset in the presence and absence of the finite-T-induced term as input data in computation of the distance responses. Table 2 provides descriptive statistics for the distance distributions based on these two statistical populations. While the sampling distributions in Figs. 3 and 4 show that both data capture critical aspects of the distance response function, Table 2 highlights the clear advantage of the 8-parameter population. It achieves a smaller average distance to the 2018 Planck result, indicating higher accuracy of the 8-parameter estimation method. Note that this superiority of the 8-parameter method is achieved despite using fewer data points, nearly 100,000 fewer than the 7-parameter competitor and relying on a coarser grid. In contrast, the 7-parameter method, with finer grid sampling and more extensive data, falls short of the prediction accuracy attained by the 8-parameter method. This numerical comparison indicates the effectiveness of our proposed 8-parameter approach, demonstrating its ability to deliver better predictions with less computational effort and fewer resources.
@@
Statistics with finite-T effects without finite-T effects
@@
Figure 5: The training and validation loss functions of the proposed ANN method in estimating the distance response function based on eight cosmological parameters (including the finite-T effects) with over 5000 epochs.
@@
In the following, we treat the output function as the distance between the estimation method and the 2018 Planck data. Hence, the superiority of the methods in predicting the Planck function is translated into the superiority of the estimation methods in minimization of the distance response function based on the data generated from the distance function under the two statistical populations. To evaluate the performance of the methods in minimizing the distance response function, we identified 20,000 data points with the smallest distance values from each statistical population. From now on, we will focus on these two data sets as our statistical populations to find the minimum of the distances. Since the functional form of the distance from these statistical populations is unknown, first, one needs to predict the distance as a function of the cosmological features in $\mathbb{R}^{7}$ and $\mathbb{R}^{8}$ corresponding to the two statistical populations in absence and presence of the finite-T terms, respectively. To do so, we applied ANNs to numerically predict the distance functions from these two statistical populations.
@@
We train the ANNs-based model, described in section 2.2.1 (see Fig. 2 for a schematic diagram of the trained ANNs), separately for the 7-parameter and 8-parameter statistical populations. For the 8-parameter population, we employ an ANN that begins with an input layer of eight nodes, one for each cosmological parameter, including the finite-T terms, followed by four hidden layers, each containing 16 neurons. We use Rectified Linear Unit (ReLU) activation functions throughout the hidden layers to obtain a robust capability for modelling complex, nonlinear relationships among the parameters in 8D. Before training, all input features are standardized to account for the broad range of magnitudes present in the different coordinates. To do so, each parameter is transformed to zero mean and unit variance. This step facilitates faster convergence and mitigates potential biases introduced by vastly different parameter scales. We split the dataset into 80% training and 20% test data sets to evaluate prediction performance to validate that model selection and hyperparameter tuning on unseen test data. The network is trained using the Adam optimizer with a learning rate of 0.001 with a mean squared error (MSE) loss function (9) over 5000 epochs. For fair comparisons, we train and compute the ANN-based prediction performance of the distance function using a 7-parameter statistics population under the same ANN configuration as described above. By combining a carefully standardized feature set, a structured architecture of hidden layers, and a well-tuned optimization, the ANNs effectively capture the dependencies underlying the distance response in both statistical populations.
@@
Figs. 5 and 6 show the training and test loss values based on 8- and 7-parameter statistical populations, respectively. It is easy to see that the training and test loss functions resemble patterns very closely, confirming the ANN-based estimates could predict the underlying distance in both statistical populations. Comparison of Figs. 5 and 6 reveals an interesting contrast. Although the eight-parameter model is constructed from a more coarsely sampled statistical population, its ANN estimator’s test losses track the training losses more closely than the seven-parameter model. This difference appears to stem from the inclusion of the finite-T density parameters, which significantly enhances the estimator’s ability to capture variability of the distance without imposing any over-fitting in the prediction problem. Consequently, the eight-parameter test loss converges faster, whereas the seven-parameter model shows a slight shortfall, evidenced by larger gaps between its training and test loss curves.

### v2

@@ 2.2 Modification of CLASS
With the conceptual framework established in Section 2.1, we now turn to the implementation of these ideas in CLASS. In its default configuration, CLASS evolves a cosmological model derived from the action
$S=S_{\mathrm{CC}}+S_{\mathrm{EH}}+S_{\mathrm{kin}},$ (18)
where $S_{\mathrm{CC}}$ denotes the cosmological constant term, $S_{\mathrm{EH}}$ is the Einstein–Hilbert action, and $S_{\mathrm{kin}}$ collectively represents the kinetic theory terms for matter and radiation. We modify CLASS to accommodate the extended system in which the standard cosmological constant contribution, denoted by $\Lambda_{1}$, is replaced by the effective quantity $\Lambda$ defined in Eq. (10).
[…]
$h,\qquad\omega_{b},\qquad\omega_{cdm},\qquad A_{s},\qquad n_{s},\qquad\tau_{reio}.$ (19)
@@
[…]
Referring to [6] and prior studies on the 1PI effective action for the Standard Model coupled to gravity, we find that the expression for the Hubble parameter $H(t)$ is modified. The original expression, as adopted in [11], is
$H$ $=$ $H_{0}\left[\Omega_{M}\Big(\frac{T}{T_{\gamma_{0}}}\Big)^{3}+\Omega_{R}\Big(\frac{T}{T_{\gamma_{0}}}\Big)^{4}\right]^{\frac{1}{2}}$ (20)
$=$ $7.204\times 10^{-19}\;T^{\frac{3}{2}}\sqrt{1.523\times 10^{-5}T+\Omega_{M}h^{2}}.$
where $T$ denotes the temperature. (In [11], both the $\Omega_{\Lambda_{1}}$ and curvature density parameter are neglected for their small contributions.) The modified CLASS implementation [26], [^8] The code was obtained in collaboration with K.-T. Cho and M.-S. Lee. includes finite-temperature corrections, yielding the following form:
$\hskip-21.68121ptH$ $=$ $7.204\times 10^{-19}\;T^{\frac{3}{2}}\sqrt{1.523\times 10^{-5}T\!+\!\left(\frac{2.725}{T}\right)^{3}\Omega_{\Lambda_{1}}h^{2}+\Omega_{M}h^{2}\!+\!\frac{T}{2.725}\Omega_{\Lambda_{2}}h^{2}\!+\!\frac{2.725}{T}\Omega_{\Lambda_{3}}h^{2}}$
. (21)
@@ 3 Numerical analysis
Our analysis employs a modified version of CLASS incorporating the quantum gravity modifications previously described. The parameter space exploration proceeds in two distinct phases. First, we conduct a comprehensive brute-force scan across eight cosmological parameters spanning generous ranges to ensure complete coverage of the parameter space. By these comprehensive scans tens of millions of data points are obtained. Following this initial broad exploration, we implement a refined search strategy focusing on regions that minimize the Euclidean distance (see the comments below) between our computed power spectra and the Planck dataset. This targeted approach yields significantly improved precision, with the minimal achieved distance of 26.83 corresponding to parameter values largely consistent with existing literature. [^9] To give a sense of the accuracy associated with this number, the distance between the 2018 Planck curve and one generated by CLASS using the default.ini settings is approximately 75. Any curve with a distance to the Planck curve of less than approximately 1000 is visually indistinguishable from it. Notably, the optimal solution features a negative value for $\Omega_{\Lambda_{2}}$, a result that finds natural explanation through dimensional regularization in the renormalization procedure.
The methodology described above [^10] While we use frequentist inference, as noted in [15], this approach deviates from traditional frequentist methods: cosmological parameters are treated as input variables in the regression framework, with the polynomial coefficients taking on the role of statistical parameters instead of the cosmological parameters themselves. The resulting model serves as a surrogate, approximating the relationship between cosmological parameters and observables. departs significantly from the standard Bayesian parameter inference employed by the Planck collaboration[18], and several comments are in order to clarify this choice. The primary motivation for adopting a non-standard, deterministic approach lies in the behavior of the newly introduced parameters, particularly $\Omega_{\Lambda_{2}}$. Although theoretical estimates for the magnitude of $\Omega_{\Lambda_{2}}$ are possible and in fact necessary as a preliminary step, it was observed in [15] that these estimates overshoot the best-fit value by several orders of magnitude (a possible explanation for this mismatch is discussed in [15]) - a discrepancy that results in power spectra incompatible with observational data. Given the lack of reliable priors and the computational expense of performing full Bayesian inference over such a vast and poorly constrained parameter space, we instead adopted an iterative trial-and-error method. This involved progressively reducing $\Omega_{\Lambda_{2}}$ and performing wide parameter scans until convergence toward the observed Planck 2018 power spectrum was achieved. Although this procedure lacks formal statistical underpinnings (see the conclusion for further discussion), it proved effective in identifying the correct order of magnitude for $\Omega_{\Lambda_{2}}$.
Our analysis leveraged the unbinned theoretical power spectrum file provided by the Planck 2018 collaboration:
COM_PowerSpect_CMB-base-plikHM-TTTEEE-lowl-lowE-lensing-minimum-theory_R3.01.txt. As this dataset represents a smooth, noise-free theoretical prediction sampled at integer multipoles and lacks an associated covariance matrix, it further justifies our departure from likelihood-based methods that rely on detailed error modeling. In contrast, the observed spectra (given, e.g., in COM_PowerSpect_CMB-TT-full_R3.01.txt) are binned and require covariance information to be properly interpreted, which was not necessary for our purposes. A major advantage of our approach is its computational efficiency. The entire parameter scan process was executed within the CLASSy framework, with only a minimal external script used to calculate and rank distance measures - an operation that completes nearly instantaneously. In comparison, integrating additional tools like Cobaya or MontePython for full Bayesian inference increased runtime by a factor of six to eight. Given that our data collection alone required over three months, the reduction in computational overhead was a decisive factor. While our trial-and-error method is not a substitute for a rigorous statistical analysis, it enables rapid, resource-efficient exploration of the parameter space. Once a viable range for $\Omega_{\Lambda_{2}}$ is identified, more focused Bayesian scans using Cobaya become both feasible and significantly less time-consuming.
@@
To complement and enhance the efficiency of our parameter space exploration, we implement advanced ANN techniques. Various numerical calculations and machine learning methods have been recently applied to the physics of black hole solutions[41, 42, 43, 40]. These machine learning methods enable precise identification of optimal parameter coordinates while maintaining computational tractability across the challenging eight-dimensional parameter space.
@@
Let us explore the estimation of the distance based on the currently established parameters as well as the finite-T induced ones, $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$. Throughout this section, the prediction accuracies are translated into the performance of the ANN-based method in detecting the minimum distance to the 2018 Planck curve from statistical populations in the presence of the finite-T effects.
@@
Figure 2: Sampling distribution of the distance values based on 8 cosmological parameters.
@@
The numerical analyses evaluate the distance between the 2018 Planck curve and 7-parameter and 8-parameter ones using 459,912 and 335,985 data points, respectively. These data points are computed over equally spaced grids spanning the domain of the two approaches. These two data sets are considered as two statistical populations corresponding to the two estimation methods. We consider the response variable $y$, representing the distance function $y=f({\bf x})$, which measures the deviation of the proposed estimation methods from the 2018 Planck power spectrum. In this setup, the parameters of the estimation methods were treated as explanatory variables ${\bf x}=(x_{1},\ldots,x_{J})$, where $J=7$ for the 7-parameter model and $J=8$ for the 8-parameter model. Using this framework, the ML model maps the cosmological parameters to the distance function, enabling us to evaluate and predict the performance of the two estimation methods. This approach was particularly valuable for assessing the accuracy of the proposed methods, in the presence or absence of $\Omega_{\Lambda_{2}}$.
@@
Figure 3: Sampling distribution of the distance values based on 7 cosmological parameters.
@@
Figs. 2 and 3 show the sampling distributions of the distance responses based on two statistical populations derived from the entire dataset in the presence and absence of the leading-order finite-T effect parameter, $\Omega_{\Lambda_{2}}$, as input data in computation of the distance responses. Table 2 provides descriptive statistics for the distance distributions based on these two statistical populations. While the sampling distributions in Figs. 2 and 3 show that both data capture critical aspects of the distance response function, Table 2 highlights the clear advantage of the 8-parameter population. It achieves a smaller average distance to the 2018 Planck result, indicating higher accuracy of the 8-parameter estimation method. Note that this superiority of the 8-parameter method is achieved despite using fewer data points, nearly 100,000 fewer than the 7-parameter competitor and relying on a coarser grid. In contrast, the 7-parameter method, with finer grid sampling and more extensive data, falls short of the prediction accuracy attained by the 8-parameter method. This numerical comparison indicates the effectiveness of our proposed 8-parameter approach, demonstrating its ability to deliver better predictions with less computational effort and fewer resources.
@@
Statistics with $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ with $\Omega_{\Lambda_{3}}$
@@
Figure 4: The training and validation loss functions of the proposed ANN method in estimating the distance response function based on eight cosmological parameters with over 5000 epochs.
@@
In the following, we treat the output function as the distance between the estimation method and the 2018 Planck data. Hence, the superiority of the methods in predicting the distance function is translated into the superiority of the estimation methods in minimization of the distance response function based on the data generated from the distance function under the two statistical populations. To evaluate the performance of the methods in minimizing the distance response function, we identified 20,000 data points with the smallest distance values from each statistical population. From now on, we will focus on these two data sets as our statistical populations to find the minimum of the distances. Since the functional form of the distance from these statistical populations is unknown, first, one needs to predict the distance as a function of the cosmological features in $\mathbb{R}^{7}$ and $\mathbb{R}^{8}$ corresponding to the two statistical populations in absence and presence of $\Omega_{\Lambda_{2}}$ term, respectively. To do so, we applied ANNs to numerically predict the distance functions from these two statistical populations.
@@
We train the ANNs-based model, described in section B.1 (see Fig. 11 for a schematic diagram of the trained ANNs), separately for the 7-parameter and 8-parameter statistical populations. For the 8-parameter population, we employ an ANN that begins with an input layer of eight nodes, one for each cosmological parameter, followed by four hidden layers, each containing 16 neurons. We use Rectified Linear Unit (ReLU) activation functions throughout the hidden layers to obtain a robust capability for modelling complex, nonlinear relationships among the parameters in 8D. Before training, all input features are standardized to account for the broad range of magnitudes present in the different coordinates. To do so, each parameter is transformed to zero mean and unit variance. This step facilitates faster convergence and mitigates potential biases introduced by vastly different parameter scales. We split the dataset into 80% training and 20% test data sets to evaluate prediction performance to validate that model selection and hyperparameter tuning on unseen test data. The network is trained using the Adam optimizer with a learning rate of 0.001 with a mean squared error (MSE) loss function (B.3) over 5000 epochs. For fair comparisons, we train and compute the ANN-based prediction performance of the distance function using a 7-parameter statistics population under the same ANN configuration as described above. By combining a carefully standardized feature set, a structured architecture of hidden layers, and a well-tuned optimization, the ANNs effectively capture the dependencies underlying the distance response in both statistical populations.
@@
Figs. 4 and 5 show the training and test loss values based on 8- and 7-parameter statistical populations, respectively. It is easy to see that the training and test loss functions resemble patterns very closely, confirming the ANN-based estimates could predict the underlying distance in both statistical populations. Comparison of Figs. 4 and 5 reveals an interesting contrast. Although the eight-parameter model is constructed from a more coarsely sampled statistical population, its ANN estimator’s test losses track the training losses more closely than the seven-parameter model. Consequently, the eight-parameter test loss converges faster, whereas the seven-parameter model shows a slight shortfall, evidenced by larger gaps between its training and test loss curves.